An 18-ft board is cut into two pieces, one of which is 4 ft longer than the other. How long is each piece?
Answer by stanbon(57361)

An 18-ft board is cut into two pieces, one of which is 4 ft longer than the other. How long is each piece?
Let the shorter piece have length "x" ft.
The longer piece is "x+4" ft. long.
EQUATION:
x+x+4=18 ft.
2x+4=18
2x=14
x=7 ft (length of the shorter piece)
x+4= 11 ft (length of the longer piece)
Cheers,
Stan H.